# A silicon spin vacuum: isotopically enriched 28silicon-on-insulator and 28silicon from ultra-high fluence ion implantation

### v1

@@ Abstract
Isotopically enriched silicon (Si) can greatly enhance qubit coherence times by minimizing naturally occurring 29Si which has a non-zero nuclear spin. Ultra-high fluence 28Si ion implantation of bulk natural Si substrates was recently demonstrated as an attractive technique to ultra-high 28Si isotopic purity. In this work, we apply this 28Si enrichment process to produce 28Si and 28Si-on-insulator (SOI) samples. Experimentally, we produced a 28Si sample on natural Si substrate with 29Si depleted to 7 ppm (limited by measurement noise floor), that is at least 100 nm thick. This is achieved with an ion energy that results in a sputter yield of less than one and a high ion fluence, as supported by simulations. Further, our simulations predict the 29Si and 30Si depletion in our sample to be less than 1 ppm. In the case of SOI, ion implantation conditions are found to be more stringent than those of bulk natural Si in terms of minimizing threading dislocations upon subsequent solid phase epitaxy annealing. Finally, we report the observation of nanoscopic voids in our 28SOI and 28Si samples located in the depth region between the surface and 70 nm. These voids appear to be stabilized by the presence of gold impurities and can be annealed out completely by a standard SPE annealing protocol at 620C for 10 minutes in the absence of gold.
@@ I Introduction
Ultra-high fluence 28Si self-implantation ($\gtrsim 10^{18}$ cm-2) can be used as a means to isotopically enrich natural silicon (natSi) substrates to unprecedented purity levels. [1, 2] With the fast growing interest in quantum technology research, 28Si substrates are important to Si-based quantum information processing (QIP) platforms such as gate-defined quantum dot qubits, [3, 4, 5] flip-flop donor qubits, [6, 7, 8] and Er-doped Si quantum emitters [9, 10] due to the zero nuclear spin of 28Si isotopes and weak spin-orbit coupling in Si. natSi consists of 4.7 at.of the non-zero nuclear spin isotope 29Si $(I=1/2)$, leading to decoherence of qubits by spectral diffusion. [11] The other stable isotope, 30Si, present to 3.1 at.has $I=0$ and does not contribute to decoherence via magnetic interactions, however, it can introduce higher order perturbations to the Hamiltonian via strain interactions. [12] An ideal substrate would therefore consist of isotopically pure monocrystalline 28Si, the much sought-after ‘semiconductor spin vacuum’. [13]
The ultra-high fluence 28Si self-implantation process, which we refer to as the ‘28Si enrichment process’, presents itself as a cost-effective and competitive alternative to existing 28SiF4-based growth techniques [14, 15, 16, 17, 18, 19, 20] and molecular beam epitaxy (MBE). [21] 28SiF4-based techniques rely on toxic and expensive isotopically pure 28SiF4 gas as its source material. The resultant Si purity is limited by that of the source and can result in purities as low as 50 ppm 29Si. This material has been used in the Avogadro project to define the kilogram. [22] A similar level of isotopic purity can also be achieved through solid source MBE (50 ppm 29Si) which uses 28Si wafers as its source material. [21] In contrast, the 28Si enrichment process employed in the present work uses natSi as both its source and starting materials. A natSi ion beam is accelerated through a standard magnetic field-based charge-to-mass analyzer and the 28Si ion beam is mass selected. Through the continuous bombardment of this 28Si ion beam into a natSi substrate, 29Si and 30Si are gradually sputtered away, resulting in isotopically purified 28Si surface layers with a thickness of $\sim 100$ nm. [1] The ion bombardment results in a transformation to the amorphous silicon phase and a subsequent moderate to high temperature anneal leads to solid phase epitaxial (SPE) regrowth of single crystal silicon with negligible diffusion of the implanted atoms expected. [23, 24, 25] A four to five-fold improvement in the 31P donor electron spin coherence time as compared to natSi has been achieved with a 0.3 at.29Si (3000 ppm) sample produced by 28Si enrichment. [1] Indeed, the donor coherence time measured in Ref. 26 was limited by inter-donor coupling/instantaneous diffusion, thus leading one to expect even longer qubit coherence times in these isotopically enriched 28Si samples. [25]
@@
Furthermore, it is well-known from both simulations and experiments that there is a spatial separation between the vacancy and interstitial distributions generated by the ion collisions during implantation. [31, 32] An excess of vacancies is expected near the surface up to approximately half the ion projected range ($R_{p}/2$) while an excess of interstitials is expected deeper in the substrate up to around one to two times the ion projected range (1 to $2\times R_{p}$). For Si implanted under non-amorphizing conditions, this spatial separation results in the formation of open volume defects (vacancy clusters, voids) and interstitial clusters (loops, rods, dislocations) in these depth ranges, respectively. [33, 34, 35, 36, 37] These defects have been observed experimentally and are stable up to temperatures of 850 to 900C. For Si implanted under amorphizing conditions, to our knowledge, there is no experimental evidence for the presence of similar complex defect clusters. [29, 38, 39, 40] Nevertheless, we expect the same spatial separation of vacancies and interstitials in Si following amorphizing ion implantation, especially in high fluence implantation such as that considered in this work. The interplay between point defect generation (determined by the ion flux), clustering (determined by the ion flux and sample temperature) and annihilation (influenced by the sample temperature) rates will therefore dictate whether open volumes can form and stabilize near the sample surface. [28] Further, these open volumes (if present) will be in close proximity to qubits (usually located within 20 nm of the surface) and thus it is important to investigate the possibility of their formation and thermal stability under further processing, as explored in this work.
In this work, we investigate the 28Si enrichment process using a range of ion beam flux and fluence values. We also consider the impact of ion implantation energy and sample temperature on both the 28Si enrichment and the crystal quality after SPE annealing. Importantly, the 28Si enrichment process is also applied to natSi-on-insulator (natSOI) substrates. SOI substrates are of particular interest for optical interfacing with Er spins [9, 10] and mechanical resonators [41, 42], and for gate-defined quantum dot architectures [3, 4] where isotopic purity may be advantageous. We find that, although the 28Si enrichment process is successfully applied to SOI, a more stringent set of ion implantation conditions must be observed to achieve successful single crystal enriched 28Si due to flux and temperature effects (collectively referred to as ‘thermal effects’ in this work). This work also leads to improved modeling of the 28Si enrichment process and to materials with $<$7 ppm 29Si, which is the detection limit of our measurement system, using a fluence under $1\times 10^{19}\;\rm cm^{-2}$.
@@ II Experimental and simulation details
Samples were characterized with cross-sectional transmission electron microscopy (TEM, FEI Tecnai F20) and selected area electron diffraction (SAED) on lamellas prepared by Ga+ milling with a focus-ion-beam (ThermoFisher Aquilos G2 cryoFIB). Prior to milling, carbon (C) and platinum (Pt) protective layers were sputter coated and electron and/or ion beam deposited.
@@ III.1 SPE annealing
Figure 1: TEM images of SOI samples implanted under high flux conditions: (a) Pristine SOI; (b) SOI sample after a 60 keV $0.77\times 10^{18}$ cm-2 implantation at 77 K and (c) after SPE annealing; (d) SOI sample after a 50 keV $0.77\times 10^{18}$ cm-2 implantation at RT and (e) after SPE annealing. TEM images are aligned along the Si-SiO2 interface and shared the same scale bar shown in (a).
@@
TEM images of selected SOI samples implanted at either 77 K or RT (both under high ion flux conditions) are shown in Fig. 1. A TEM image of pristine SOI is shown for reference [Fig. 1(a)]. Erosion or accumulation caused during the enrichment process are characterized with a change in thickness with reference to this pristine sample, $\Delta z$, where $\Delta z<0$ indicates erosion, and $\Delta z>0$ accumulation. All samples in Fig. 1 exhibit accumulation with $\Delta z=16\pm 3$ and $\Delta z=9\pm 3$ nm for the samples shown Fig. 1(b) and (d), respectively. The $\Delta z$ of the corresponding samples after SPE annealing [Fig. 1(c) and (e)] are $\Delta z=15\pm 3$ and $\Delta z=9\pm 3$. We note that the density of amorphous Si is 1.8% less than that of crystalline Si [46] and therefore $\Delta z$ in our crystalline Si samples should be 2 to 4 nm less than the corresponding amorphous Si samples. This, however, was not observed in our TEM presumably due to small variations in the Si layer thickness across the SOI substrate as indicated by the supplier. The surface sputter yield, $SY$, quantifies the number of sputtered Si atoms per incident ion and can be calculated assuming negligible changes to the Si atomic density using the equation $SY=1-\frac{\Delta z}{200}\frac{10^{18}}{\Phi}$, where $\Phi$ refers to the implanted 28Si ion fluence in units of cm-2 and $\Delta z$ has units of nm.  [^1] Derived assuming a Si atomic density of $5\times 10^{22}$ atoms/cm3 and when the sputter yield is zero, $\Delta z=200$ nm for an ion fluence of $10^{18}$ cm-2. We will discuss the ion flux, fluence and energy dependencies of $\Delta z$ and $SY$ later in Sec. III.2.
@@
In Fig. 1(b), it is clear that the ultra-high fluence 28Si ion implantation (50 keV $0.77\times 10^{18}$ cm-2, high flux, 77 K) resulted in the formation of a continuous amorphous layer that extends from the surface to almost all the way through the entire Si device layer (labeled ‘a-28Si’). The amorphous nature of this implanted layer is confirmed by SAED measurements (diffuse rings, not shown). Importantly, a 20 nm thick crystalline seed underneath this amorphous layer is retained (labeled ‘c-natSi’). During SPE annealing, this crystal silicon provides a seed for the complete recrystallization of the amorphous layer, as confirmed by TEM [Fig. 1(c)] and SAED measurements (clear bright diffraction spots, not shown). A thin dark band of end-of-range (EOR) interstitial-rich defect complexes is also observed at the same depth location as the original amorphous/crystalline-interface (a/c-interface) [see Figs. 1(b-c)]. Upon SPE annealing at 620C, these interstitials accumulate to form larger and more stable defect complexes/clusters (Ostwald ripening) which can further evolve into loops and rods upon higher temperature annealing at or above 900C. [23] These rods and loop-type defects are visible at the EOR region of our samples under higher magnification TEM [data shown in Supplementary Material]. Other than this region, the recrystallized layer appears to be free from microscopic defects (to within the resolution of our TEM). However, we will demonstrate that defects may in fact remain in the form of nanoscopic voids within this amorphized 28Si layer using gold decoration experiments below.
@@
In Fig. 1(d), we show the TEM image of a RT implanted SOI sample (50 keV $0.77\times 10^{18}$ cm-2, high flux). As compared with the 77 K implanted sample shown previously in Fig. 1(b), the RT implantation also resulted in the formation of a continuous amorphous layer but with a few notable differences. First, the location of the a/c-interface is shallower in the sample shown in Fig. 1(d) as compared to Fig. 1(b) due to a lower ion implantation energy (50 keV and 60 keV, respectively). In general, for sub-100 keV implantation, the location of the a/c-interface is located at a depth of around $2\times R_{p}$ from the initial surface ($\Delta z=0$), where $R_{p}$ is the projected ion range calculated from SRIM and is proportional to the ion energy. For example, for $E_{\rm{ion}}=60$ keV, $R_{p}$ = 87 nm (observed a/c = $(222\pm 5)$ nm) while for $E_{\rm{ion}}=50$ keV, $R_{p}$ = 74 nm (observed a/c = $(179\pm 15)$ nm). Second, there is a 20 nm thick dark band of defects located at the broadened a/c-interface in the RT implanted sample (Fig. 1(d)). This is in contrast to the 77 K sample which has a much sharper a/c-interface [Fig. 1(b)]. This dark band of defects may be a result of the accumulation and clustering of a high density of Si interstitials at the a/c-interface during implantation [30] and/or dynamic annealing at room temperature which broadens the transition between the amorphous and crystalline phase, as will be discussed later. The TEM image of the same sample after the SPE annealing is shown in Fig. 1(e). Here, threading dislocations were observed to emanate from this dark defect band all the way towards the surface. This was confirmed by TEM where there is evidence for strain. In addition, SAED measurements in the crystalline 28Si layer show clear bright diffraction spots (data not shown).
@@ III.2 Surface sputtering
In Fig. 2(a), we show TRIDYN simulations of $\Delta z$ as a function of ion fluence fitted to experimental data from this work and plotted on a log scale. The experimentally measured values for $\Delta z$ were obtained through TEM analysis of various SOI samples implanted under different implantation temperatures (RT or 77 K), ion flux (high/HI or low/LO) and energies (50 keV or 60 keV), as described earlier. We note that it was not possible to determine $\Delta z$ for bulk Si substrates from TEM analysis due to the lack of a buried oxide reference but this should be possible through other surface profiling techniques not performed in this work. As for the simulations, the surface binding energy, $E_{sb}$, which is the surface potential that an atom has to overcome to escape the surface, was chosen as the fitting parameter. The simulations shown in Fig. 2(a) were calculated assuming $E_{sb}=5.17$ eV and a two-layered Si/SiO2 stack. We note that, for 50 keV and 60 keV ions, changing the substrate definition to a single layer of Si increased $SY$ by less than 1%. This corresponds to a decrease in $\Delta z$, which depends linearly on the ion fluence [see Fig. 2], by up to 6% for fluences up to $10^{19}$ cm-2. The bulk binding energy, $E_{bulk}$, is set to the recommended value of 0 eV in our simulations, [45] noting that setting $E_{bulk}=0.1$ eV had a similar effect as increasing $E_{sb}$ by 2 %. The displacement threshold energy, $E_{dsp}$, had negligible effect on the simulations performed at 50 keV and 60 keV.
@@
Vacuum contamination effects are far less significant under high flux conditions and a reasonable agreement between experiment and simulations is achieved by setting $E_{sb}=5.17$ eV as shown in Fig. 2(a). Nevertheless, we believe that the need to increase $E_{sb}$ by 10% from its default value to fit our experimental data is mainly due to low levels of residual gas contamination. Other possible explanations for the need to adjust $E_{sb}$ to fit our experiments (high flux) include modifications to $E_{sb}$ due to surface damage during implantation, [52, 53] higher than expected sample temperatures due to beam heating [26] and a poor thermal conduction between the SOI substrate and sample stage, as well as void swelling effects that we will discuss later.
@@
For a 60 keV $0.77\times 10^{19}$ cm-2 implantation, TRIDYN predicts a 29Si depletion of less than 1 ppm, in good agreement with our (noise floor limited) SIMS measurements shown in Fig. 3(b). Additional simulations reveal that 29Si depletion levels of less than 1 ppm can only be achieved when the ion energy results in surface accumulation. As shown in Fig. 2(c), surface accumulation in our experiments occurs when the ion energy is less than 3 keV or higher than 40 keV. Under these ion energy conditions, the 29Si depletion in the sample decreases rapidly as a function of ion fluence and reaches a value of about 1 ppm when the ion fluence is around $0.6\times 10^{19}$ to $2.3\times 19^{19}$ cm-2 [details in Supplementary Material]. Furthermore, by increasing the ion energy from 40 keV to 80 keV, the ion fluence required to achieve 29Si depletion levels of 1 ppm or less can be lowered from $(2.3\pm 0.4)\times 10^{19}$ cm-2 to $(5.7\pm 0.3)\times 10^{18}$ cm-2. These ion implantation conditions can be easily and realistically achieved as shown by our experiments, demonstrating the competitiveness and attractiveness of our 28Si enrichment process for producing Si spin vacuum substrates required by QIP. In the next section [Sec.III.4], we show that the isotopically enriched 28Si sub-surface layer contains nanoscopic voids that can be easily removed under moderate temperature annealing. The observation of these open volumes was enabled by the gettering of fast diffusing impurities such as gold and copper to the defects during thermal processing, [37] where the fast diffuser was deliberately introduced into the sample surface prior to thermal annealing.

### v2

@@ Abstract
Isotopically enriched silicon (Si) can greatly enhance qubit coherence times by minimizing naturally occurring 29Si which has a non-zero nuclear spin. Ultra-high fluence 28Si ion implantation of bulk natural Si substrates was recently demonstrated as an attractive technique to ultra-high 28Si isotopic purity. In this work, we apply this 28Si enrichment process to produce 28Si and 28Si-on-insulator (SOI) samples. Experimentally, we produced a 28Si sample on natural Si substrate with 29Si depleted to 7 ppm (limited by measurement noise floor), that is at least 100 nm thick. This is achieved with an ion energy that results in a sputter yield of less than one and an ultra-high ion fluence, as supported by our improved computational model that is based on fitting a large number of experiments. Further, our model predicts the 29Si and 30Si depletion in our sample to be less than 1 ppm. In the case of SOI, ion implantation conditions are found to be more stringent than those of bulk natural Si in terms of minimizing threading dislocations upon subsequent solid phase epitaxy annealing. Finally, we do not observe open volume defects in our 28SOI and 28Si samples after SPE annealing (620C, 10 minutes).
@@ I Introduction
Ultra-high fluence 28Si self-implantation ($\gtrsim 10^{18}$ cm-2) can be used as a means to isotopically enrich natural silicon (natSi) substrates to unprecedented purity levels. [1, 2] With the fast growing interest in quantum technology research, 28Si substrates are important to Si-based quantum information processing (QIP) platforms such as gate-defined quantum dot qubits, [3, 4, 5] flip-flop donor qubits, [6, 7, 8] and Er-doped Si quantum emitters [9, 10, 11] due to the zero nuclear spin of 28Si isotopes and weak spin-orbit coupling in Si. natSi consists of 4.7 at.of the non-zero nuclear spin isotope 29Si $(I=1/2)$, leading to decoherence of qubits by spectral diffusion. [12] The other stable isotope, 30Si, present to 3.1 at.has $I=0$ and does not contribute to decoherence via magnetic interactions, however, it can introduce higher order perturbations to the Hamiltonian via strain interactions. [13] An ideal substrate would therefore consist of isotopically pure monocrystalline 28Si, the much sought-after ‘semiconductor spin vacuum’. [14]
The ultra-high fluence 28Si self-implantation process, which we refer to as the ‘28Si enrichment process’, presents itself as a cost-effective and competitive alternative to existing 28SiF4-based growth techniques [15, 16, 17, 18, 19, 20, 21] and molecular beam epitaxy (MBE). [22] 28SiF4-based techniques rely on toxic and expensive isotopically pure 28SiF4 gas as its source material. The resultant Si purity is limited by that of the source and can result in purities as low as 50 ppm 29Si. This material has been used in the Avogadro project to define the kilogram. [23] A similar level of isotopic purity can also be achieved through solid source MBE (50 ppm 29Si) which uses 28Si wafers as its source material. [22] In contrast, the 28Si enrichment process employed in the present work uses natSi as both its source and starting materials. A natSi ion beam is accelerated through a standard magnetic field-based charge-to-mass analyzer and the 28Si ion beam is mass selected. Through the continuous bombardment of this 28Si ion beam into a natSi substrate, 29Si and 30Si are gradually sputtered away, resulting in isotopically purified 28Si surface layers with a thickness of $\sim 100$ nm. [1] The ion bombardment results in a transformation to the amorphous silicon phase and a subsequent moderate to high temperature anneal leads to solid phase epitaxial (SPE) regrowth of single crystal silicon with negligible diffusion of the implanted atoms expected. [24, 25, 26] A four to five-fold improvement in the 31P donor electron spin coherence time as compared to natSi has been achieved with a 0.3 at.29Si (3000 ppm) sample produced by 28Si enrichment. [1] Indeed, the donor coherence time measured in Ref. 27 was limited by inter-donor coupling/instantaneous diffusion, thus leading one to expect even longer qubit coherence times in these isotopically enriched 28Si samples. [26]
@@
Open volume defects (vacancy clusters, voids) can form in the near-surface region of crystalline Si under non-amorphizing implantation conditions. [32, 33, 34, 35, 36] This is in part due to the spatial separation between the vacancy and interstitial distributions generated by the ion collisions during implantation. [37, 38] An excess of vacancies is expected near the surface up to approximately half the ion projected range ($R_{p}/2$) while an excess of interstitials is expected deeper in the substrate up to around one to two times the ion projected range ($R_{p}$ to $2R_{p}$).
For Si implanted under amorphizing conditions, including the 28Si enrichment process, to our knowledge, there is no direct experimental evidence for the presence of these open volume defects. [1, 30, 39] This may be due to their apparent instability in amorphous Si, [40, 41] and resolution limitations of conventional microscopy techniques. The latter may be circumvented through decoration of the open volumes with trace metallic impurities to aid their observation, which may complicate subsequent interpretation. Nevertheless, these open volume defects (if present) will be in close proximity to qubits (usually located within 20 nm of the surface). Thus, it is important to consider the possibility of their formation under the 28Si enrichment process and their thermal stability under further processing, as explored in this work.
In this work, we investigate the 28Si enrichment process using a range of ion beam flux and fluence values. We also consider the impact of ion implantation energy and sample temperature on both the 28Si enrichment and the crystal quality after SPE annealing. Importantly, the 28Si enrichment process is also applied to natSi-on-insulator (natSOI) substrates. SOI substrates are of particular interest for optical interfacing with Er spins [9, 10, 11] and mechanical resonators [42, 43], and for gate-defined quantum dot architectures [3, 4] where isotopic purity may be advantageous. We find that, although the 28Si enrichment process is successfully applied to SOI, a more stringent set of ion implantation conditions must be observed to achieve successful single crystal enriched 28Si due to flux and temperature effects (collectively referred to as ‘thermal effects’ in this work). This work also leads to an improved computational model of the 28Si enrichment process, which ultimately enabled us to produce 28Si samples with $<$7 ppm 29Si, which is the detection limit of our measurement system. Our model predicts the 29Si and 30Si depletion in these samples to be less than 1 ppm.
@@ II Experimental and simulation details
Samples were characterized with cross-sectional transmission electron microscopy (TEM, FEI Tecnai F20) and selected area electron diffraction (SAED) on lamellas prepared by Ga+ milling with a focus-ion-beam (ThermoFisher Aquilos 2 cryoFIB). Prior to milling, carbon (C) and platinum (Pt) protective layers were sputter coated and electron and/or ion beam deposited.
@@ III.1 SPE annealing
Figure 1: TEM images of SOI samples implanted under high flux conditions: (a) Pristine SOI; (b) SOI sample after a 60 keV $0.77\times 10^{18}$ cm-2 implantation at 77 K and (c) after SPE annealing; (d) SOI sample after a 50 keV $0.77\times 10^{18}$ cm-2 implantation at RT and (e) after SPE annealing. TEM images are aligned along the Si-SiO2 interface and shared the same scale bar shown in (a). (f-i) SAED images obtained from the dotted regions indicated in the TEM images shown in (b-e).
@@
TEM images and SAED patterns of selected SOI samples implanted at either 77 K or RT (both under high ion flux conditions) are shown in Fig. 1. A TEM image of pristine SOI is shown for reference [Fig. 1(a)]. Erosion or accumulation caused during the enrichment process are characterized with a change in thickness with reference to this pristine sample, $\Delta z$, where $\Delta z<0$ indicates erosion, and $\Delta z>0$ accumulation. All samples in Fig. 1 exhibit accumulation with $\Delta z=16\pm 3$ and $\Delta z=9\pm 3$ nm for the samples shown Fig. 1(b) and (d), respectively. The $\Delta z$ of the corresponding samples after SPE annealing [Fig. 1(c) and (e)] are $\Delta z=15\pm 3$ and $\Delta z=9\pm 3$. We note that the density of amorphous Si is 1.8% less than that of crystalline Si [47] and therefore $\Delta z$ in our crystalline Si samples should be 2 to 4 nm less than the corresponding amorphous Si samples. This, however, was not observed in our TEM presumably due to small variations in the Si layer thickness across the SOI substrate as indicated by the supplier. The surface sputter yield, $SY$, quantifies the number of sputtered Si atoms per incident ion and can be calculated assuming negligible changes to the Si atomic density using the equation $SY=1-\frac{\Delta z}{200}\frac{10^{18}}{\Phi}$, where $\Phi$ refers to the implanted 28Si ion fluence in units of cm-2 and $\Delta z$ has units of nm.  [^1] Derived assuming a Si atomic density of $5\times 10^{22}$ atoms/cm3 and when the sputter yield is zero, $\Delta z=200$ nm for an ion fluence of $10^{18}$ cm-2. We will discuss the ion flux, fluence and energy dependencies of $\Delta z$ and $SY$ later in Sec. III.2.
@@
In Fig. 1(b), it is clear that the ultra-high fluence 28Si ion implantation (50 keV $0.77\times 10^{18}$ cm-2, high flux, 77 K) resulted in the formation of a continuous amorphous layer that extends from the surface to almost all the way through the entire Si device layer (labeled ‘a-28Si’). The amorphous nature of this implanted layer is confirmed by SAED measurements (diffuse rings, Fig. 1(f)). Importantly, a 20 nm thick crystalline seed underneath this amorphous layer is retained (labeled ‘c-natSi’). During SPE annealing, this crystal silicon provides a seed for the complete recrystallization of the amorphous layer, as confirmed by TEM [Fig. 1(c)] and SAED measurements (clear bright diffraction spots, Fig. 1(g)). A thin dark band of end-of-range (EOR) interstitial-rich defect complexes is also observed at the same depth location as the original amorphous/crystalline-interface (a/c-interface) [see Figs. 1(b-c)]. Upon SPE annealing at 620C, these interstitials accumulate to form larger and more stable defect complexes/clusters (Ostwald ripening) which can further evolve into loops and rods upon higher temperature annealing at or above 900C. [24] These rods and loop-type defects are visible at the EOR region of our samples under higher magnification TEM [data shown in Supplementary Material]. Other than this region, the recrystallized layer appears to be free from microscopic defects (to within the resolution of our TEM).
@@
In Fig. 1(d), we show the TEM image of a RT implanted SOI sample (50 keV $0.77\times 10^{18}$ cm-2, high flux). As compared with the 77 K implanted sample shown previously in Fig. 1(b), the RT implantation also resulted in the formation of a continuous amorphous layer but with a few notable differences. First, the location of the a/c-interface is shallower in the sample shown in Fig. 1(d) as compared to Fig. 1(b) due to a lower ion implantation energy (50 keV and 60 keV, respectively). In general, for sub-100 keV implantation, the location of the a/c-interface is located at a depth of around $2\times R_{p}$ from the initial surface ($\Delta z=0$), where $R_{p}$ is the projected ion range calculated from SRIM and is proportional to the ion energy. For example, for $E_{\rm{ion}}=60$ keV, $R_{p}$ = 87 nm (observed a/c = $(222\pm 5)$ nm) while for $E_{\rm{ion}}=50$ keV, $R_{p}$ = 74 nm (observed a/c = $(179\pm 15)$ nm). Second, there is a 20 nm thick dark band of defects located at the broadened a/c-interface in the RT implanted sample (Fig. 1(d)). This is in contrast to the 77 K sample which has a much sharper a/c-interface [Fig. 1(b)]. This dark band of defects may be a result of the accumulation and clustering of a high density of Si interstitials at the a/c-interface during implantation [31] and/or dynamic annealing at room temperature which broadens the transition between the amorphous and crystalline phase, as will be discussed later. The TEM image of the same sample after the SPE annealing is shown in Fig. 1(e). Here, threading dislocations were observed to emanate from this dark defect band all the way towards the surface. This was confirmed by TEM where there is evidence for strain. In addition, SAED measurements in the crystalline 28Si layer show clear bright diffraction spots, as shown in Fig. 1(i).
@@ III.2 Surface sputtering
In Fig. 2(a), we show TRIDYN simulations of $\Delta z$ as a function of ion fluence fitted to experimental data from this work and plotted on a log scale. The experimentally measured values for $\Delta z$ were obtained through TEM analysis of various SOI samples implanted under different implantation temperatures (RT or 77 K), ion flux (high/HI or low/LO) and energies (50 keV or 60 keV), as described earlier. We note that it was not possible to determine $\Delta z$ for bulk Si substrates from TEM analysis due to the lack of a buried oxide reference but this should be possible through other surface profiling techniques not performed in this work. As for the simulations, the surface binding energy, $E_{sb}$, which is the surface potential that an atom has to overcome to escape the surface, was chosen as the fitting parameter. The simulations shown in Fig. 2(a) were calculated assuming $E_{sb}=5.17$ eV and a two-layered Si/SiO2 stack. We note that, for 50 keV and 60 keV ions, changing the substrate definition to a single layer of Si increased $SY$ by less than 1%. This corresponds to a decrease in $\Delta z$, which depends linearly on the ion fluence [see Fig. 2], by up to 6% for fluences up to $10^{19}$ cm-2. The bulk binding energy, $E_{bulk}$, is set to the recommended value of 0 eV in our simulations, [46] noting that setting $E_{bulk}=0.1$ eV had a similar effect as increasing $E_{sb}$ by 2 %. The displacement threshold energy, $E_{dsp}$, had a negligible effect on the simulations performed at 50 keV and 60 keV.
@@
Vacuum contamination effects are far less significant under high flux conditions and a reasonable agreement between experiment and simulations is achieved by setting $E_{sb}=5.17$ eV as shown in Fig. 2(a). We believe that the need to increase $E_{sb}$ by 10% from its default value to fit our experimental data is primarily due to low levels of residual gas contamination. In contrast, the implantation temperature and other thermal effects, including local beam heating [27] and poor thermal conduction to the sample stage, are expected to decrease $E_{sb}$ by less than $2\%$. This estimate is obtained by assuming that our samples have reached temperatures close to 500–600C during implantation, in which case the thermal energy is $\approx 1.5\%$ of $E_{sb}$. Note that it is very unlikely for our samples to have exceeded 600C given the amorphous nature of the implanted layer, as shown previously in Fig. 1, and that the thermal energy near the Si melting point (1400C, $k_{B}T=0.15$ eV) is only $3\%$ of $E_{sb}$. Similarly, increased surface damage due to higher ion fluxes can also be expected to decrease $E_{sb}$. [53, 54] Further experiments under high quality vacuum ($<5\times 10^{-8}$ Torr) over a wider range of ion fluxes are required to quantify its reducing effect on $E_{sb}$.
@@
For a 60 keV $0.77\times 10^{19}$ cm-2 implantation, TRIDYN predicts a 29Si depletion of less than 1 ppm, in good agreement with our (noise floor limited) SIMS measurements shown in Fig. 3(b). Additional simulations reveal that 29Si depletion levels of less than 1 ppm can only be achieved when the ion energy results in surface accumulation. As shown in Fig. 2(c), surface accumulation in our experiments occurs when the ion energy is less than 3 keV or higher than 40 keV. Under these ion energy conditions, the 29Si depletion in the sample decreases rapidly as a function of ion fluence and reaches a value of about 1 ppm when the ion fluence is around $0.6\times 10^{19}$ to $2.3\times 19^{19}$ cm-2 [details in Supplementary Material]. Furthermore, by increasing the ion energy from 40 keV to 80 keV, the ion fluence required to achieve 29Si depletion levels of 1 ppm or less can be lowered from $(2.3\pm 0.4)\times 10^{19}$ cm-2 to $(5.7\pm 0.3)\times 10^{18}$ cm-2. These ion implantation conditions can be easily and realistically achieved as shown by our experiments, demonstrating the competitiveness and attractiveness of our 28Si enrichment process for producing Si spin vacuum substrates required by QIP. In the next section [Sec.III.4], we show through gold decoration experiments that the isotopically enriched 28Si samples are free from nanoscopic voids after SPE annealing.